# Let Me Speak Freely? A Study on the Impact of Format Restrictions on Performance of Large Language Models

### v2

@@ Abstract
Zhi Rui Tam1thanks: Equal contribution, Cheng-Kuang Wu [^1] footnotemark: 1, Yi-Lin Tsai1, Chieh-Yen Lin1, Hung-yi Lee2thanks: Equal advisorship, Yun-Nung Chen [^2] footnotemark: 2 1Appier AI Research 2National Taiwan University
@@ 3.3 Evaluation method
Perfect Text Parser. To disentangle format errors from the actual performance of the generated content, we use an LLM prompted to extract the final answer from the text, rather than relying on regex or string parsers. This approach acts as a perfect parser, minimizing errors introduced when switching between different models. Our ablation study, comparing different models, found that claude-3-haiku-20240307 is the most consistent when using gpt-4-turbo as a human reference, compared to four other low-cost APIs. Full results can be found in Appendix B.
@@ 5.2 Comparison Across Different Formats
On hindsight we do not see any structure format which consistency stands out from others which generalized across all models in Figure 4. For Gemini model, we found JSON is more consistent however it does not always outperform other format.
@@ 5.3 Structure Format and Parsing Error Rates
Table 2: Parsing error percentage across different models
@@ Appendix E Comparison between using regex and LLM as answer parser in GSM8K
Table 5: Alignment between GPT-4-turbo and human-parsed answers
@@ F.1 Zero shot prompting comparing Text, JSON, XML, YAML
Table (8, 7) shows all the number with standard deviation on all 4 format (NL, JSON, XML, YAML) in classification and reasoning tasks.

### v3

@@ Abstract
Zhi Rui Tam1thanks: Equal contribution , Cheng-Kuang Wu [^1] footnotemark: 1, Yi-Lin Tsai1, Chieh-Yen Lin1, Hung-yi Lee2, Yun-Nung Chen2 1Appier AI Research, 2National Taiwan University
@@ 1 Introduction
  • We explore not only JSON but also other commonly used schemas like XML and YAML. Additionally, we test three different format-restricting strategies: constrained decoding, format-restricting instructions, and NL-to-Format, all of which are applicable to industrial settings.
@@ 3.1.2 Classification Tasks
3.2 Output Format
When designing the output format for each format, we wish to keep the schema simple; hence, we limit the number of key-value pairs for each dataset to 2: reasoning and answer fields. On top of this limitation, we permute the naming of the field names (e.g., "reasoning", "step-by-step reasoning").
While the outputs in our study may appear simplistic, converting Large Language Model (LLM) responses to a desired format is not trivial in practice. LLMs’ output often deviates from instructions, necessitating complex parsing code to handle various response variations and edge cases, particularly when separating reasoning from the final answer. This problem is exacerbated when switching between different LLMs, as each model may have its own preferred output format, potentially breaking existing parser code. We have encountered this issue numerous times when building LLM applications, often resorting to instructing LLMs to respond in structured formats (e.g., JSON) to reduce the complexity of our parser code.
Our choice of simple output structures (one reasoning and one final answer field) was deliberate, allowing us to focus on the impact of structural bias on LLM reasoning ability, which is the primary aim of our work. We acknowledge that exploring LLM robustness with more complex output structures would be valuable. We have noted this as an important direction for future research.
@@ 3.4 Evaluation method
Perfect Text Parser. To disentangle format errors from the actual performance of the generated content, we use an LLM prompted to extract the final answer from the text, rather than relying on regex or string parsers. This approach acts as a perfect parser, minimizing errors introduced when switching between different models. Our ablation study, comparing different models, found that claude-3-haiku-20240307 is the most consistent when using gpt-4-turbo as a human reference, compared to four other low-cost APIs. Detailed comparison between gpt-4-turbo between human parsed answers as well as comparison of other LLMs can be found in Appendix B.
@@ 5.2 Comparison Across Different Formats
On hindsight we do not see any structure format which consistency stands out from others which generalized across all models in Figure 4. For Gemini model, we found JSON is more consistent however it does not always outperform other format for example Claude-3-Haiku.
@@ 5.3 Structure Format and Parsing Error Rates
5.4 Study on Structure Generation with Context-free Grammars
A newer revision of the model gpt-4o-mini-2024-07-18 now supports Context-free Grammars via a so-called Structure Output API. This API allows users to provide a predefined JSON schema, ensuring the response adheres to it with 100% guarantee. It’s important to note that this differs from the previously mentioned JSON-mode on OpenAI’s models, which uses the OpenAI function calling API. We conducted experiments on 3 reasoning datasets using gpt-4o-mini, denoting the newer structured output method as JSON-schema. Results are shown in Table 2.
Task NL FRI JSON-Mode JSON-Schema
GSM8K 94.57 87.17 86.95 91.71
(3.95) (4.43) (1.36) (0.68)
Shuffle Obj 82.85 81.46 76.43 81.77
(5.67) (3.71) (9.74) (6.86)
Last Letter 83.11 84.73 76.00 86.07
(3.54) (2.99) (6.69) (3.33)
Table 2: Performance of gpt-4o-mini-2024-07-18 across tasks and formats. In 2 out of 3 reasoning datasets, NL (Natural Language) still performs slightly better than JSON-Schema.
@@ 5.4 Study on Structure Generation with Context-free Grammars
Table 3: Parsing error percentage across different models. We want to highlight that despite having near zero parsing error in Gemini-Flash XML and YAML, there’s still degradation in the final benchmark scores.
@@ Appendix B Choosing which LLMs as answer extraction
We first validate if existing LLMs such as gpt-4-turbo can the perfect parser in answer extraction in reasoning tasks such as GSM8K, Last Letter Concatenation. We sampled 300 responses in total: 100 each from Last Letter, Shuffle Object, and GSM8K, each of the responses were independently parsed by human evaluators. We then compared the human-parsed answers with those extracted by GPT-4-turbo. The result shown in Table 4, shows gpt-4-turbo can indeed denote as a perfect parser in these 3 cases.
Task Examples Accuracy (%)
Last Letter 100 97.0
Shuffle Obj 100 96.0
GSM8K 100 100.0
Average 300 97.7
Table 4: Alignment between GPT-4-Turbo and human annotations across different tasks.
@@ Appendix E Comparison between using regex and LLM as answer parser in GSM8K
Table 7: Alignment between GPT-4-turbo and human-parsed answers. In general we found GPT-4-turbo is very close to perfect parser which serves as a versatile parser to all kinds of task.
@@ F.1 Zero shot prompting comparing Text, JSON, XML, YAML
Table (10, 9) shows all the number with standard deviation on all 4 format (NL, JSON, XML, YAML) in classification and reasoning tasks.